# OAT-FM: Optimal Acceleration Transport for Improved Flow Matching

### v1

@@ Abstract
Figure 1: (a) The principle of OAT-FM and the corresponding two-phase FM paradigm. (b) We train SiT-XL (Ma et al., 2024) by our two-phase FM paradigm on the ImageNet $\times256$ dataset (Deng et al., 2009).
@@ 1 Introduction
As a promising generative modeling strategy, Flow Matching (FM) (Song et al., 2021a; Ho et al., 2020; Lipman et al., 2023) aims to learn a (deterministic or stochastic) velocity field capturing the transport of probability mass from a prior noise distribution to a complex data distribution. In general, the learned velocity field corresponds to a neural solver of a specific ordinary or stochastic differential equation. Therefore, in the inference phase, we can simulate the differential equation with discrete sampling steps, resulting in the flows from noise to data. Nowadays, flow matching has achieved competitive performance in various challenging generative tasks, e.g., image generation (Lipman et al., 2023; Esser et al., 2024), audio generation (Liu et al., 2024; Wang et al., 2024), protein design (Bose et al., 2024; Yue et al., 2025), and so on.
@@
Currently, some two-phase FM paradigms are proposed to achieve high-quality generation efficiently. Typically, given a well-trained flow/diffusion model, Rectified Flow (ReFlow) (Liu et al., 2023) and its variants (Lee et al., 2024; Hu et al., 2025) iteratively refit each learned flow trajectory to its linear or piecewise linear approximation. Such a so-called rectification phase straightens the flow progressively and thus allows a large sampling step size (and thus requires few steps) during inference. In addition, self-distillation methods, like Consistency Distillation (CD) (Song et al., 2023; Yang et al., 2024), treat the given model as a teacher and learn an efficient student model to fit the flow trajectories created by the teacher. Essentially, both these two-phase FM paradigms utilize the velocity field of the given model in their phase-2 training, enhancing model efficiency while preventing severe performance degradation.
@@
In this study, we propose a novel improved FM method, called OAT-FM, based on the recent theory of Optimal Acceleration Transport (OAT) (Chen et al., 2018; Brigati et al., 2025), which provides a new way to leverage velocity information and leads to a new two-phase FM paradigm. As illustrated in Figure 1(a), OAT-FM minimizes the acceleration transport between the noise distribution and the data one defined in the product space of samples and their velocities. The implementation of OAT-FM corresponds to a bi-level optimization problem. In the lower-level, we solve an OAT problem, leveraging the endpoint velocity of flow to compute the Optimal Transport (OT) plan (or called coupling (Villani, 2021)) defined in the product space. In the context of FM, we can decompose the coupling and solve the OAT problem efficiently, whose complexity is the same with that of the classic OT problem (Peyré and Cuturi, 2019). Accordingly, we can sample noise-data pairs based on the OT plan during training. In the upper-level, we minimize an upper bound for the optimal acceleration transport from noise to data, which straightens flow trajectories with a theoretical guarantee. As shown in Figure 1(a), given an arbitrary flow/diffusion model, whose velocity information has been relatively reliable, we can continually refine it via our OAT-FM method and improve its performance.
@@
Different from existing OT-based FM methods (e.g., OT-CFM (Tong et al., 2024; Pooladian et al., 2023), OFM (Kornilov et al., 2024), and ReFlow (Liu, 2022; Hertrich et al., 2025)) and recent acceleration-driven FM methods (Chen et al., 2025a, b; Cao et al., 2025; Gong et al., 2025), OAT-FM applies a physically grounded objective tied directly to acceleration control and transport, whose straightening objective is mathematically equivalent to minimizing the physical action associated with acceleration defined by OAT. Compared with existing two-phase FM paradigms, such as ReFlow and CD, the paradigm based on OAT-FM neither requires generating a large number of paired training data nor relies on dense intermediate interpolation results, which eliminates the risk of data distribution drift. To our knowledge, OAT-FM makes the first attempt to explore the usefulness of acceleration-driven FM in high-dimensional generative tasks. As shown in Figure 1(b), the phase-2 training achieved by OAT-FM leads to promising high-resolution image generation results.
@@ 2.1 Optimal Transport-based Flow Matching
Most existing FM methods fall into a generalized Conditional Flow Matching (CFM) framework (Tong et al., 2024). Denote $\mathbb{P}(\mathcal{X})$ as the set of probability measures defined in a sample space $\mathcal{X}$. Suppose that we have a data distribution $\rho_{1}\in\mathbb{P}(\mathcal{X})$ and a noise one $\rho_{0}\in\mathbb{P}(\mathcal{X})$, respectively. Typically, we set $\rho_{0}$ as a normal distribution $\mathcal{N}(0,1)$. CFM models the evolutionary sample distribution from $\rho_{0}$ to $\rho_{1}$ conditioned on an auxiliary variable $z$, called conditional path and denoted as $p_{t}(x|z)$ (with $t\in[0,1]$). It learns a neural network, denoted as $v_{\theta}$, to fit the velocity field $v_{t}(x|z)$ corresponding to $p_{t}(x|z)$, i.e.,
@@
In the CFM framework, the distribution $\pi$ and the conditional path $p_{t}(x|z)$ play a central role, and different implementations result in various FM methods (Lim et al., 2024; Tong et al., 2024). For example, the early FM method in (Lipman et al., 2023) sets $\pi=\rho_{1}$ and $p_{t}(x|z)$ as a Gaussian distribution $\mathcal{N}\big(tz,(t\sigma-t+1)^{2}\big)$. The independent coupling CFM (I-CFM) sets $\pi=\rho_{0}\times\rho_{1}$, i.e., $z=(x_{0},x_{1})$ with $x_{0}\sim\rho_{0}$ and $x_{1}\sim\rho_{1}$ independently. Recently, some attempts have been made to interpret FM through the lens of optimal transport (Villani, 2021), leading to a series of optimal transport-based FM methods, e.g., OT-CFM (Tong et al., 2024; Pooladian et al., 2023) OFM (Kornilov et al., 2024), and kinetic FM (Shaul et al., 2023, 2025).
Optimal Transport-based CFM (OT-CFM): Denote the Wasserstein-2 distance between $\rho_{0}$ and $\rho_{1}$ as $\mathcal{W}_{2}(\rho_{0},\rho_{1})$. The dynamic (Benamou–Brenier) formulation of optimal transport (Benamou and Brenier, 2000) seeks a unique least–kinetic–energy flow $v$ corresponding to $\mathcal{W}_{2}^{2}(\rho_{0},\rho_{1})$, i.e.,
@@
In theory, the objective function in (3) regresses $v_{\theta}(x_{t},t)$ to the constant velocity $x_{1}-x_{0}$. This constant velocity equals the characteristic velocity on the Wasserstein geodesic when $(x_{0},x_{1})$ are coupled optimally (McCann, 1997; Dong et al., 2024). In other words, training with $\pi^{\ast}$ aligns $v_{\theta}(x,t)$ with the least–kinetic–energy flow in (2), yielding straighter and more efficient flow trajectories. However, constant velocity is sufficient but not necessary for straightening flows.
@@
Motivated by Proposition 1, we can move beyond first-order dynamics and minimize acceleration instead, which corresponds to the optimal acceleration transport problem in (Chen et al., 2018; Brigati et al., 2025) and leads to the proposed OAT-FM method accordingly.
@@ 2.2 Flow Matching Based on Optimal Acceleration Transport
Definition 1 (Dynamic Formulation of Optimal Acceleration Transport (OAT)).
@@
subject to the Vlasov equation (Vlasov, 1968) $\partial_{t}\mu+v\cdot\nabla_{x}\mu+\nabla_{v}\cdot\big(a\,\mu\big)=0$, with boundary conditions $\mu(\cdot,\cdot,0)=\mu_{0}$ and $\mu(\cdot,\cdot,1)=\mu_{1}$. Here, $a:\mathcal{X}\times\mathcal{V}\times[0,1]\mapsto\mathbb{R}^{d}$ is the acceleration field, and the Vlasov equation expresses conservation of mass in the product space.
Similar to the first-order optimal transport problem, OAT admits a static coupling problem on the product space in the Kantorovich format.
Definition 2 (Kantorovich formulation of OAT (Brigati et al., 2025)).
@@
The learning problem in (7) considers an OAT-based flow-matching objective. In particular, the parameter $\alpha$ balances directional alignment with a proxy for small total acceleration, and the expectation over $t$ averages these effects along the path. The objective of OAT-FM provides a tight bound on the true OAT second-order discrepancy, ensuring effective minimization of the acceleration.
@@ 2.3 Efficient Implementation
Similar to OT-CFM, the OAT-FM problem in (7) is a bi-level optimization problem as well. The lower-level problem determines a coupling $\pi^{\ast}$. The upper-level problem updates $\theta$ given $\pi^{\ast}$. In practice, we solve them via alternating optimization. Following the existing methods in (Pooladian et al., 2023; Tong et al., 2024), we solve the lower-level problem via min-batch approximation.
It should be noted that, although the coupling of OAT problem has a four-dimensional coupling, i.e., $\pi(z_{0},z_{1})=\pi(x_{0},x_{1},v_{0},v_{1})$, it has a decomposable structure in the context of FM. In particular, given an arbitrary sample $x$ and a time stamp $t$, we can determine its velocity as $v_{\theta}(x,t)$, which is conditionally independent of other samples or velocities. Therefore, we have $\pi(x_{0},x_{1},v_{0},v_{1})=\pi_{x}(x_{0},x_{1})\pi(v_{0},v_{1}|x_{0},x_{1})=\pi_{x}(x_{0},x_{1})\pi_{v}(v_{0}|x_{0})\pi_{v}(v_{1}|x_{1})$, where $\pi_{x}$ is the marginal coupling associated with sample pairs and $\pi_{v}(\cdot|x_{t})=\delta_{v_{\theta}(x,t)}$. As a result, the lower-level OAT problem in (7) can be reformulated as
$\mathop{\arg\min}_{\pi\in\Pi(\mu_{0},\mu_{1})}\mathbb{E}_{(z_{0},z_{1})\sim\pi}\Big[c_{\mathcal{A}}^{2}(z_{0},\,z_{1})\Big]\Rightarrow\mathop{\arg\min}_{\pi_{x}\in\Pi(\rho_{0},\rho_{1})}\mathbb{E}_{(x_{0},x_{1})\sim\pi_{x}}\Big[\|x_{1}-x_{0}\|^{2}-(x_{1}-x_{0})^{\top}\bar{v}\Big],$ (9)
where $\rho_{0},\rho_{1}\in\mathbb{P}(\mathcal{X})$ denote the noise and data distributions, and $\bar{v}=\mathbb{E}_{v_{0}\sim\pi_{v}(\cdot|x_{0}),v_{1}\sim\pi_{v}(\cdot|x_{1})}[v_{1}+v_{0}]$. The reformulated problem in (9) becomes a classic OT problem (Peyré and Cuturi, 2019). As a result, the complexity of OAT-FM is the same as OT-CFM (Tong et al., 2024). The detailed derivation of (9) and the scheme of our learning algorithm are provided in Appendix B.
@@ 3 Comparisons with Related Work
Different from existing OT-based flow matching methods (Tong et al., 2024; Pooladian et al., 2023; Kornilov et al., 2024), OAT-FM considers the coupling in the “sample-velocity” product space that minimizes the expected acceleration of flow, which fully leverages the endpoint velocity of flow to compute the OT plan and sampling noise-data pairs. Alongside OAT-FM, several works have extended diffusion/flow models based on second-order dynamics. Let $\rho_{t}\in\mathbb{P}(\mathcal{X})$. For $x_{0}\sim\rho_{0}$ and $x_{1}\sim\rho_{1}$, the rectified interpolation (Gong et al., 2025, Definition 3.12) between them, including the trajectory, the velocity, and the acceleration, can be written as $x_{t}=\alpha_{t}x_{0}+\beta_{t}x_{1}$, $v_{t}=\dot{\alpha}_{t}x_{0}+\dot{\beta}_{t}x_{1}$, and $a_{t}=\ddot{\alpha}_{t}x_{0}+\ddot{\beta}_{t}x_{1}$, where $\{\alpha_{t},\beta_{t},\dot{\alpha}_{t},\dot{\beta}_{t},\ddot{\alpha}_{t},\ddot{\beta}_{t}\}$ are predefined time–dependent coefficients. When $\rho_{t}$ satisfies the continuity equation $\partial_{t}\rho_{t}(x)+\nabla_{x}\cdot\big(v_{t}(x),\rho_{t}(x)\big)=0$, differentiating it once more in time yields a second–order continuity law (Gong et al., 2025, Lemma 5.1):
@@
Building on this perspective, recent second–order flow matching methods incorporate acceleration into their training losses. NRFlow (Chen et al., 2025b) augments the standard OT-CFM objectives by regressing a neural velocity predictor $v_{\theta_{1}}(x,t)$ to the rectified velocity $v_{t}$, and simultaneously regressing an acceleration predictor $a_{\theta_{2}}(v,x,t)$ to the rectified acceleration $a_{t}$. The resulting loss is
@@
HOMO (Chen et al., 2025a) extends NRFlow by adding a self-consistency penalty, ensuring that the instantaneous velocity matches the average of the velocity and its one-step update $x_{t+d}$, i.e.,
@@
with the target velocity defined as $\bar{v}_{t}=\tfrac{1}{2}(v_{\theta_{1}}(x_{t},t)+v_{\theta_{1}}(x_{t+d},t+d))$. Inspired by MeanFlow (Geng et al., 2025), SOM (Cao et al., 2025) replaces pointwise supervision with interval averages and aligns the model to averaged quantities over $[r,t]$, whose training loss is

### v2

@@ Abstract
Figure 1: (a) The principle of OAT-FM and the corresponding two-phase FM paradigm. (b) We train SiT-XL (ma2024sit) by our two-phase FM paradigm on the ImageNet $\times256$ dataset (deng2009imagenet).
@@ 1 Introduction
As a promising generative modeling strategy, Flow Matching (FM) (Song2021ScoreBased; Ho2020Denoising; lipman2023flow) aims to learn a (deterministic or stochastic) velocity field capturing the transport of probability mass from a prior noise distribution to a complex data distribution. In general, the learned velocity field corresponds to a neural solver of a specific ordinary or stochastic differential equation. Therefore, in the inference phase, we can simulate the differential equation with discrete sampling steps, resulting in the flows from noise to data. Nowadays, flow matching has achieved competitive performance in various challenging generative tasks, e.g., image generation (lipman2023flow; esser2024scaling), audio generation (liu2024generative; wang2024frieren), protein design (bose2024se; yue2025reqflow), and so on.
@@
Currently, some two-phase FM paradigms are proposed to achieve high-quality generation efficiently. Typically, given a well-trained flow/diffusion model, Rectified Flow (ReFlow) (Liu2022RectifiedFlow) and its variants (lee2024improving; hu2025improving) iteratively refit each learned flow trajectory to its linear or piecewise linear approximation. Such a so-called rectification phase straightens the flow progressively and thus allows a large sampling step size (and thus requires few steps) during inference. In addition, self-distillation methods, like Consistency Distillation (CD) (song2023consistency; yang2024consistency), treat the given model as a teacher and learn an efficient student model to fit the flow trajectories created by the teacher. Essentially, both these two-phase FM paradigms utilize the velocity field of the given model in their phase-2 training, enhancing model efficiency while preventing severe performance degradation.
@@
In this study, we propose a novel improved FM method, called OAT-FM, based on the recent theory of Optimal Acceleration Transport (OAT) (Chen2018MeasureValued; Brigati2025KineticOT), which provides a new way to leverage velocity information and leads to a new two-phase FM paradigm. As illustrated in Figure 1(a), OAT-FM minimizes the acceleration transport between the noise distribution and the data one defined in the product space of samples and their velocities. The implementation of OAT-FM corresponds to a bi-level optimization problem. In the lower-level, we solve an OAT problem, leveraging the endpoint velocity of flow to compute the Optimal Transport (OT) plan (or called coupling (villani2021topics)) defined in the product space. In the context of FM, we can decompose the coupling and solve the OAT problem efficiently, whose complexity is the same with that of the classic OT problem (peyre2019computational). Accordingly, we can sample noise-data pairs based on the OT plan during training. In the upper-level, we minimize an upper bound for the optimal acceleration transport from noise to data, which straightens flow trajectories with a theoretical guarantee. As shown in Figure 1(a), given an arbitrary flow/diffusion model, whose velocity information has been relatively reliable, we can continually refine it via our OAT-FM method and improve its performance.
@@
Different from existing OT-based FM methods (e.g., OT-CFM (tong2024improving; pooladian2023multisample), OFM (kornilov2024optimal), and ReFlow (liu2022rectified; hertrich2025relation)) and recent acceleration-driven FM methods (chen2025high; chen2025nrflow; cao2025towards; gong2025theoretical), OAT-FM applies a physically grounded objective tied directly to acceleration control and transport, whose straightening objective is mathematically equivalent to minimizing the physical action associated with acceleration defined by OAT. Compared with existing two-phase FM paradigms, such as ReFlow and CD, the paradigm based on OAT-FM neither requires generating a large number of paired training data nor relies on dense intermediate interpolation results, which eliminates the risk of data distribution drift. To our knowledge, OAT-FM makes the first attempt to explore the usefulness of acceleration-driven FM in high-dimensional generative tasks. As shown in Figure 1(b), the phase-2 training achieved by OAT-FM leads to promising high-resolution image generation results.
@@ 2.1 Optimal Transport-based Flow Matching
Most existing FM methods fall into a generalized Conditional Flow Matching (CFM) framework (tong2024improving). Denote $\mathbb{P}(\mathcal{X})$ as the set of probability measures defined in a sample space $\mathcal{X}$. Suppose that we have a data distribution $\rho_{1}\in\mathbb{P}(\mathcal{X})$ and a noise one $\rho_{0}\in\mathbb{P}(\mathcal{X})$, respectively. Typically, we set $\rho_{0}$ as a normal distribution $\mathcal{N}(0,1)$. CFM models the evolutionary sample distribution from $\rho_{0}$ to $\rho_{1}$ conditioned on an auxiliary variable $z$, called conditional path and denoted as $p_{t}(x|z)$ (with $t\in[0,1]$). It learns a neural network, denoted as $v_{\theta}$, to fit the velocity field $v_{t}(x|z)$ corresponding to $p_{t}(x|z)$, i.e.,
@@
In the CFM framework, the distribution $\pi$ and the conditional path $p_{t}(x|z)$ play a central role, and different implementations result in various FM methods (lim2024elucidating; tong2024improving). For example, the early FM method in (lipman2023flow) sets $\pi=\rho_{1}$ and $p_{t}(x|z)$ as a Gaussian distribution $\mathcal{N}\big(tz,(t\sigma-t+1)^{2}\big)$. The independent coupling CFM (I-CFM) sets $\pi=\rho_{0}\times\rho_{1}$, i.e., $z=(x_{0},x_{1})$ with $x_{0}\sim\rho_{0}$ and $x_{1}\sim\rho_{1}$ independently. Recently, some attempts have been made to interpret FM through the lens of optimal transport (villani2021topics), leading to a series of optimal transport-based FM methods, e.g., OT-CFM (tong2024improving; pooladian2023multisample) OFM (kornilov2024optimal), and kinetic FM (shaul2023kinetic; shaul2025flow).
Optimal Transport-based CFM (OT-CFM): Denote the Wasserstein-2 distance between $\rho_{0}$ and $\rho_{1}$ as $\mathcal{W}_{2}(\rho_{0},\rho_{1})$. The dynamic (Benamou–Brenier) formulation of optimal transport (BenamouBrenier2000) seeks a unique least–kinetic–energy flow $v$ corresponding to $\mathcal{W}_{2}^{2}(\rho_{0},\rho_{1})$, i.e.,
@@
In theory, the objective function in (3) regresses $v_{\theta}(x_{t},t)$ to the constant velocity $x_{1}-x_{0}$. This constant velocity equals the characteristic velocity on the Wasserstein geodesic when $(x_{0},x_{1})$ are coupled optimally (Mccann1997ConstantVelocity; dong2024monge). In other words, training with $\pi^{\ast}$ aligns $v_{\theta}(x,t)$ with the least–kinetic–energy flow in (2), yielding straighter and more efficient flow trajectories. However, constant velocity is sufficient but not necessary for straightening flows.
@@
Motivated by Proposition 1, we can move beyond first-order dynamics and minimize acceleration instead, which corresponds to the optimal acceleration transport problem in (Chen2018MeasureValued; Brigati2025KineticOT) and leads to the proposed OAT-FM method accordingly.
@@ 2.2 Flow Matching Based on Optimal Acceleration Transport
Definition 1 (Dynamic Formulation of Optimal Acceleration Transport (OAT) (benamou2019second)).
@@
subject to the Vlasov equation (vlasov1968vibrational) $\partial_{t}\mu+v\cdot\nabla_{x}\mu+\nabla_{v}\cdot\big(a\,\mu\big)=0$, with boundary conditions $\mu(\cdot,\cdot,0)=\mu_{0}$ and $\mu(\cdot,\cdot,1)=\mu_{1}$. Here, $a:\mathcal{X}\times\mathcal{V}\times[0,1]\mapsto\mathbb{R}^{d}$ is the acceleration field, and the Vlasov equation expresses conservation of mass in the product space.
Similar to the first-order optimal transport problem, OAT admits a static coupling problem on the product space in the Kantorovich format.  [^2] To our knowledge, the Kantorovich formulation of OAT is first proposed and discussed by (Chen2018MeasureValued, Eq.10), and see also (benamou2019second, Eq.4.14) and (Brigati2025KineticOT, Eq.8).
Definition 2 (Kantorovich formulation of OAT (Chen2018MeasureValued; benamou2019second; Brigati2025KineticOT)).
@@
The learning problem in (7) considers an OAT-based flow-matching objective. In particular, the parameter $\alpha$ balances directional alignment with a proxy for small total acceleration, and the expectation over $t$ averages these effects along the path.
Remark. The connection between the OAT problem in (5) and our OAT-FM method is analogous to that between the OT problem and OT-CFM. In particular, it has been well known that OT-CFM learns a flow to achieve an optimal transport in the sample space, straightening flow trajectories by pursuing constant velocity (tong2024improving). Our OAT-FM learns a flow to achieve optimal acceleration transport in the product space of sample and velocity, straightening flow trajectories by minimizing acceleration (i.e., a smooth velocity field).
The objective of OAT-FM provides a tight bound on the true OAT second-order discrepancy, ensuring effective minimization of the acceleration.
@@ 2.3 Efficient Implementation
Similar to OT-CFM, the OAT-FM problem in (7) is a bi-level optimization problem as well. The lower-level problem determines a coupling $\pi^{\ast}$. The upper-level problem updates $\theta$ given $\pi^{\ast}$. In practice, we solve them via alternating optimization. Following the existing methods in (pooladian2023multisample; tong2024improving), we solve the lower-level problem via min-batch approximation.
It should be noted that, although the coupling of OAT problem has a four-dimensional coupling, i.e., $\pi(z_{0},z_{1})=\pi(x_{0},x_{1},v_{0},v_{1})$, it has a decomposable structure in the context of FM. In particular, given an arbitrary sample $x$ and a time stamp $t$, we can determine its velocity as $v_{\theta}(x,t)$, which is conditionally independent of other samples or velocities. Therefore, we have $\pi(x_{0},x_{1},v_{0},v_{1})=\pi_{x}(x_{0},x_{1})\pi(v_{0},v_{1}|x_{0},x_{1})=\pi_{x}(x_{0},x_{1})\pi_{v}(v_{0}|x_{0})\pi_{v}(v_{1}|x_{1})$, where $\pi_{x}$ is the marginal coupling associated with sample pairs and $\pi_{v}(\cdot|x_{t})=\delta_{v_{\theta}(x,t)}$. As a result, we simplify the lower-level OAT problem in (7) as
$\mathop{\arg\min}_{\pi\in\Pi(\mu_{0},\mu_{1})}\!\!\mathbb{E}_{(z_{0},z_{1})\sim\pi}\Big[c_{\mathcal{A}}^{2}(z_{0},z_{1})\Big]\!\!\Rightarrow\!\!\mathop{\arg\min}_{\pi_{x}\in\Pi(\rho_{0},\rho_{1})}\!\!\mathbb{E}_{(x_{0},x_{1})\sim\pi_{x}}\Big[12\|x_{1}\!-\!x_{0}-\bar{v}_{x_{0},x_{1}}\|^{2}\!+\!\|\tilde{v}_{x_{0},x_{1}}\|_{2}^{2}\Big],\!\!$ (9)
where $\rho_{0},\rho_{1}\in\mathbb{P}(\mathcal{X})$ denote the noise and data distributions, $\bar{v}_{x_{0},x_{1}}=\frac{1}{2}\big(v_{\theta}(x_{0},0)+v_{\theta}(x_{1},1)\big)$, and $\tilde{v}_{x_{0},x_{1}}=v_{\theta}(x_{1},1)-v_{\theta}(x_{0},0)$. The reformulated problem in (9) becomes a classic OT problem (peyre2019computational). As a result, the complexity of OAT-FM is the same as OT-CFM (tong2024improving). The detailed derivation of (9) and the scheme of our learning algorithm are provided in Appendix B.
@@ 3 Comparisons with Related Work
Different from existing OT-based flow matching methods (tong2024improving; pooladian2023multisample; kornilov2024optimal), OAT-FM considers the coupling in the “sample-velocity” product space that minimizes the expected acceleration of flow, which fully leverages the endpoint velocity of flow to compute the OT plan and sampling noise-data pairs. Alongside OAT-FM, several works have extended diffusion/flow models based on second-order dynamics. Let $\rho_{t}\in\mathbb{P}(\mathcal{X})$. For $x_{0}\sim\rho_{0}$ and $x_{1}\sim\rho_{1}$, the rectified interpolation (gong2025theoretical, Definition 3.12) between them, including the trajectory, the velocity, and the acceleration, can be written as $x_{t}=\alpha_{t}x_{0}+\beta_{t}x_{1}$, $v_{t}=\dot{\alpha}_{t}x_{0}+\dot{\beta}_{t}x_{1}$, and $a_{t}=\ddot{\alpha}_{t}x_{0}+\ddot{\beta}_{t}x_{1}$, where $\{\alpha_{t},\beta_{t},\dot{\alpha}_{t},\dot{\beta}_{t},\ddot{\alpha}_{t},\ddot{\beta}_{t}\}$ are predefined time–dependent coefficients. When $\rho_{t}$ satisfies the continuity equation $\partial_{t}\rho_{t}(x)+\nabla_{x}\cdot\big(v_{t}(x),\rho_{t}(x)\big)=0$, differentiating it once more in time yields a second–order continuity law (gong2025theoretical, Lemma 5.1):
@@
Building on this perspective, recent second–order flow matching methods incorporate acceleration into their training losses. NRFlow (chen2025nrflow) augments the standard OT-CFM objectives by regressing a neural velocity predictor $v_{\theta_{1}}(x,t)$ to the rectified velocity $v_{t}$, and simultaneously regressing an acceleration predictor $a_{\theta_{2}}(v,x,t)$ to the rectified acceleration $a_{t}$. The resulting loss is
@@
HOMO (chen2025high) extends NRFlow by adding a self-consistency penalty, ensuring that the instantaneous velocity matches the average of the velocity and its one-step update $x_{t+d}$, i.e.,
@@
with the target velocity defined as $\bar{v}_{t}=\tfrac{1}{2}(v_{\theta_{1}}(x_{t},t)+v_{\theta_{1}}(x_{t+d},t+d))$. Inspired by MeanFlow (geng2025mean), SOM (cao2025towards) replaces pointwise supervision with interval averages and aligns the model to averaged quantities over $[r,t]$, whose training loss is